11. The temperature coefficient of resistance for carbon at 0oC is
-0.0048/oC. What is the significance of the minus sign?
A carbon resistor has a resistance of 500Ω.Determine its resistance at 50oC
1
Expert's answer
2020-05-04T12:43:51-0400

1. The significance of the minus sign is the following: resistance of carbon decreases with increasing temperature, as one can conclude from the following formula:

"R(T) = R(T_0)(1 + \alpha\Delta T)" ,

where "R(T)" is the resistence at temperature T, "R(T_0)" is the resistance at temperature "T_0" (in our case "T_0=0\degree C" ), "\Delta T =T-T_0" and "\alpha" is the temperature coefficient of resistance.


2. From the previous equation:

"R(T) = 500\Omega(1 -0.0048\frac{1}{\degree C}(50\degree C-0\degree C)) = 380\Omega".